OpenStudy (anonymous):

I have 3 questions I will fan medal and give you a testimonial if you can help me. The data in the table are linear use the table to find the slope 3/2 - 3/2 - 2/3 2/3

OpenStudy (anonymous):

find the slope of a line that is perpendicular to the line containing the points (-2,-1) and (2, -3) m = -2 m = 2 m = -1/2 m = -1

OpenStudy (anonymous):

the ans for this question is m=2

OpenStudy (anonymous):

actually i didn't read your previous question carefully and andwered . Well the ans for your next Ques was m=2

OpenStudy (anonymous):

as you said line contains the points (-2,-1), (2,-3). So the slope of this line is [-3-(-1)]/2-(-2) which is equal to -1/2. so slope of this line is -1/2 and slope of line perpendicular to i can be found by formula m1*m2=-1 that is -1/2*m=-1 which gives the slope of perpendicular line as 2

OpenStudy (loser66):

to me, m1 = 1/m2 only, so it should be -2. but if you think you are right, I have nothing to say

OpenStudy (anonymous):

uhh

OpenStudy (loser66):

let me recheck, dont' worry, if I am wrong, I will confess that I am wrong. quite fair at that field

OpenStudy (anonymous):

ok

OpenStudy (loser66):

yes, he is right, my bad, sorry, it's =2

OpenStudy (anonymous):

doesn't matter

OpenStudy (anonymous):

ok thx everyone